the electron of hydrogen atom rotates in a circular orbit with radius 0.53 A with velocity 2.2*10 power 6 m/s Determine the intensity occurred given charge of electron = 1.6*10 power -19
1
Expert's answer
2019-07-17T11:15:49-0400

Let's calculate what we can calculate, since "the intensity" is a word with broad meaning and we need to know the intensity of what is to be calculated.

According to classical explanation, when an electron circles the nucleus, it emits light of frequency


"f=\frac{1}{T}=\frac{v}{2\pi r}=\frac{2.2\cdot10^6}{2\cdot3.142\cdot0.53\cdot10^{-10}}=6.61\cdot10^{15}\text{ Hz}."

Thus the electron should gradually lose its energy and fall on the nucleus.